# SCSA: Exploring the Synergistic Effects Between Spatial and Channel Attention

### v1

@@ III-C Synergistic Effects
TABLE III: Ablation studies on the design strategy of SCSA, conducted at a 224x224 resolution, using the ImageNet-1K validation set. The abbreviation ”PC” denotes progressive compression. $Gi(K_{1},K_{2},\ldots,K_{i})$ denotes splitting $X$ into $i$ sub-features and applying a 1D convolution of size $K_{i}$ to each $i$-th sub-feature.
@@
Baseline SCSA-50 2009 77.49
Macro Design w/o SMSA 2217 76.72
@@
TABLE IV: Comparison of our proposed SCSA with other state-of-the-art attention mechanisms across multiple benchmark models at a 224x224 resolution on the ImageNet-1K validation set.
@@
ResNet-50 ResNet[32] 25.56 4.02 2433 76.39 93.09
Channel ECANet[5] 25.56 4.11 2109 77.05 93.43
SENet[1] 28.07 4.11 2077 77.23 93.56
FcaNet[3] 28.07 4.11 1905 77.29 93.64
Hybrid CPCA[19] 27.40 4.87 1379 75.80 92.57
CBAM[11] 28.07 4.12 1687 77.12 93.50
SANet[16] 25.56 4.12 1493 77.12 93.64
ELA[18] 25.59 4.11 2233 77.25 93.52
CA[15] 25.69 4.11 2244 77.37 93.52
EMA[17] 25.57 4.18 1861 77.43 93.79
@@
ResNet-101 ResNet[32] 44.55 7.83 1588 77.76 93.81
Channel ECANet[5] 44.55 7.83 1408 78.32 93.99
SENet[1] 49.30 7.84 1399 78.40 94.05
FcaNet[3] 49.29 7.84 1242 78.51 94.10
Hybrid CBAM[11] 49.30 7.84 1118 78.09 94.07
CA[15] 44.80 7.84 1437 78.11 93.92
@@
MobileNetV2-1.0 MobileNetV2[33] 3.51 0.31 6693 71.54 90.11
Channel ECANet[5] 3.51 0.31 5746 72.02 90.35
Hybrid CBAM[11] 4.07 0.32 4539 72.43 90.49
@@
TABLE V: Comparison of the performance of different attention mechanisms for object detection on MSCOCO val2017, utilizing models such as Faster R-CNN, Cascade R-CNN, and RetinaNet. All models were fine-tuned using the ”1×1$\times$1 ×” schedule.
@@
IV Experiments
@@
In this section, we first introduce the experimental details. Next, we conduct experiments on four visual tasks, comparing our proposed SCSA with other state-of-the-art attention mechanisms. Following this, in Section IV-E, we perform a comprehensive ablation study on our meticulously designed SCSA from four different perspectives.
IV-A Implementation Details
@@
To evaluate our proposed SCSA on ImageNet-1K[35], we select three mainstream backbone networks based on CNN architectures, including ResNet-50, ResNet-101[32] and MobileNetV2-1.0[33]. Specifically, for models based on ResNets, we employ an SGD optimizer with a momentum of 0.9, a weight decay of 1e-4, and an initial learning rate of 0.05. which is reduced tenfold every 30 epochs. The models are trained using a batch size of 128 over 100 training epochs. For training MobileNetV2 with our SCSA, we follow the settings used in ECANet[5], the optimizer has a momentum of 0.9 and a weight decay of 4e-5, with an initial learning rate of 0.045. This rate decrease linearly by a factor of 0.98, and the batch size is set to 96 for 400 training epochs. Notably, to enhance training efficiency, we employ Automatically Mixed Precision(AMP) training on a single NVIDIA RTX 4090 GPU for the classification tasks.
@@
We evaluate our SCSA on MSCOCO 2017[36] using Faster R-CNN[37], Mask R-CNN[38], Cascade R-CNN[39], and RetinaNet[40]. These detectors are implemented using the MMDetection[41] toolboxes with default settings. Input images are scaled proportionally by their shorter side to 800. All models are trained using an SGD optimizer with a momentum of 0.9 and a weight decay of 1e-4, with a batch size of 2 per GPU, over a total of 12 epochs. Faster R-CNN, Mask R-CNN and Cascade R-CNN started with a learning rate of 0.0025, while RetinaNet starts at 0.00125. The learning rates for all models are decreased by a factor of 10 at the 8th and 11th epochs. We fine-tuned the model on the MSCOCO train2017 dataset for 12 epochs using a single NVIDIA H800 GPU and reported comparative results on val2017.
@@
We further validate our method on ADE20K[42] with the UperNet[43] for semantic segmentation. Following common practices[44, 45], we also utilize the MMSegmentation[46] toolboxes, set the batch size to 16, and conduct 80k training iterations. All models are trained using an SGD optimizer with an initial learning rate of 0.01, a momentum of 0.9, and a weight decay of 5e-4. We also conduct training and inference using a single NVIDIA H800 GPU.
@@ IV-B Image Classification
We compare our SCSA against other state-of-the-art attention mechanisms such as SENet[1], CBAM[11], ECANet[5], FcaNet[3], CA[15], SANet[16], EMA[17], CPCA[19], and ELA[18]. As shown in Figure 1 and Table IV, our SCSA achieved the highest Top-1 accuracy across networks of different scales, with negligible parameter count and computational complexity. Within hybrid architectures, our method’s inference speed based on ResNet is second only to CA, but it offer a better balance of accuracy, speed, and model complexity with a moderate model width.
IV-C Object Detection
We evaluate the generalization ability of our SCSA on general detection tasks to verify its effectiveness in enhancing feature extraction. We use ResNet-50 and ResNet-101 as the backbone networks and FPN[47] as the feature fusion sub-network.
As shown in Table V, our method outperforms other state-of-the-art attention methods across various detectors, model sizes, and object scales. For Faster R-CNN, our SCSA improves by 1.7% and 1.3% compared to the original ResNet-50 and ResNet-101, respectively. We also provide visualizations based on MSCOCO val2017 in Appendix C and detailed experiments on more complex scenarios (e.g., small targets on the VisDrone2019[48] dataset, dark environments on the ExDark[49] dataset, and infrared scenes on the FLIR-ADAS-V2[50] dataset) in Appendix E. These further demonstrating the effectiveness and generalizability of our method.
IV-D Segmentation
We also test its performance in semantic segmentation on ADE20K and instance segmentation on MSCOCO 2017. Some methods (e.g., CBAM, CPCA, SE, SA, ECA) have not previously been tested on ADE20K for semantic segmentation, so we conduct extensive comparative experiments based on the UperNet[43] network. As shown in Tables I and II, our SCSA significantly outperforms other attention methods. Specifically, SCSA improves performance by 0.94% and 1.02% on ResNet-50 and ResNet-101, respectively, while other methods only achieve improvements of 0.1% to 0.2%, and some even fall below the baseline model. These results demonstrate that our method, based on multi-semantic spatial information, performs exceptionally well in pixel-level tasks.
IV-E Ablation Study
As shown in Table III, we use SCSA-50 as a baseline on ImageNet-1K for ablations across four aspects.
IV-E1 Macro Design
We separately validate the SMSA and PCSA modules. Both demonstrate significant accuracy improvements, with SMSA notably enhancing classification accuracy by 1.05%. Without progressive compression in PCSA, accuracy drops by 0.18%, and this is primarily because the channel attention mechanism cannot utilize the discriminative spatial priors provided by SMSA for its computations. Additionally, when replacing the single attention head in PCSA with multi-head and channel shuffle operation, performance decrease from 77.49% to 77.35%. This phenomenon is primarily attributed to the strong inter-channel interactions facilitated by the single head, which effectively alleviate semantic disparities observed in SMSA.
IV-E2 Ordering
When PCSA is moved ahead of SMSA, there is a 0.29% drop in Top-1 accuracy. This decrease is attributed to the early channel attention, where spatial information across features remains unmodulated by spatial attention, lacking precise spatial priors to guide channel recalibration. This effectively validates our hypothesis that spatial attention can guide channel learning. Concerning the timing of normalization, placing GN before the attention computation leads to a slight reduction in accuracy. This could be due to pre-normalization diminishing the distinct semantic patterns inherent to different sub-features, thus impacting the representation of diverse semantic features.
IV-E3 Micro Design
Replacing GN with BN results in a decrease in both accuracy and inference speed, with the Top-1 accuracy dropping from 77.49% to 77.19%. This decline is attributed to GN’s superior ability to preserve the independence of semantic patterns among sub-features, thereby minimizing semantic interference. Conversely, BN is highly sensitive to batch size, introducing additional batch statistical noise for multi-semantic information. These insights suggest that GN may be a more suitable choice in convolution layers that involve multiple semantics. Furthermore, the decline in accuracy and increase in parameters with unshared convolutions further validate the effectiveness of using shared convolutions to consistently learn and model features and dependencies across the H and W dimensions.
IV-E4 Branch
Reducing the diversity of convolution kernels led to a decrease in accuracy, underscoring the significant impact of multi-semantic information on enhancing feature extraction capabilities. Drawing on the synergistic approach proposed in our SCSA, we leverage semantic space information to bolster feature extraction and employ channel self-attention to facilitate semantic interaction, effectively mitigating disparities across different semantic spaces and promoting better fusion of multi-semantic information.
V Analysis
V-A Visualization of Attention
As shown in Figure 3, our SCSA distinctly focuses on multiple key regions under similar receptive field conditions, significantly minimizing critical information loss, and providing rich feature information for the ultimate downstream tasks. This advantage arises from the SCSA’s synergistic design, which preserves critical information in both spatial and channel domains during their attention computations, further emphasizing its superior representational capabilities compared to other attention mechanisms. Detailed analysis and visualization results for detection and segmentation tasks can be found in Appendix.
VI Conclusion
In this paper, we explore the synergy between spatial and channel dimensions through dimension decoupling, lightweight multi-semantic guidance, and semantic disparities mitigation, proposing SCSA—a novel, plug-and-play spatial and channel synergistic attention mechanism. Extensive experiments across multiple visual tasks demonstrate the enhanced performance and robust generalization capabilities of our SCSA compared to other state-of-the-art attention mechanism. We hope that our research will facilitate the exploration of synergistic properties across multiple dimensions in various domains.
VII Acknowledgments
This work was supported by the National Natural Science Foundation of China (62376252); Key Project of Natural Science Foundation of Zhejiang Province (LZ22F030003).
Appendix A Integrating SCSA into the Block
Figure 4 shows our SCSA module integrated into blocks of ResNet and MobileNetV2. In our work, for ResNet, we consider the appropriate model width, which differs slightly from other attention mechanisms. For MobileNetV2, we account for the DWConv’s need to compensate for its lack of expressive capability by increasing the model width. Therefore, we place it in the same position as other attention mechanisms.
Figure 4: Main Module Structures with SCSA: (a) MobileNetv2’s inverted residual module. (b) Residual blocks for ResNet-18 and ResNet-34. (c) Residual blocks for ResNet-50 and above.
Appendix B Visualization of Segmentation Results
B-A Instance Segmentation
To validate the effectiveness of incorporating synergistic multi-semantic information into SCSA, we select three random samples from MSCOCO val2017 for instance segmentation with Mask R-CNN. As shown in Figure 5, our method segments obscured and overlapping objects more comprehensively and accurately, achieving higher confidence scores. These results underscore the benefits of our method in leveraging multi-semantic information to better perceive the contextual space of relevant objects.
Figure 5: Visualization of instance segmentation results using the Mask R-CNN detector. Each instance depicted in a distinct color.
B-B Semantic Segmentation
In addition to instance segmentation, we visualize semantic segmentation outcomes as well. Utilizing the ADE20K dataset, we randomly select three samples for inference via UperNet and compared our method visually with the ResNet-50 baseline. It can be observed from Figure 6 that our method significantly improves the segmentation of objects that overlap and are semantically adjacent, effectively distinguishing between scenarios such as spectators seated on chairs and toilets near bathtubs.
Figure 6: Visualization of semantic segmentation results using the UperNet model.
Figure 7: Detection results are visualized on Faster R-CNN, Cascade R-CNN, and RetinaNet by respectively selecting two random samples from COCO val2017 and comparing our SCSA with the ResNet-50 baseline to demonstrate the effectiveness of our method.
TABLE VI: Comparison of our SCSA, based on ResNet-50 and ResNet-101, with other attention mechanisms for object detection performance across four different datasets.
@@ B-B Semantic Segmentation
TABLE VII: Code and datasets assets used in our work.
Name URL
ImageNet-1K https://image-net.org/
MSCOCO 2017 https://cocodataset.org/
ADE20K http://sceneparsing.csail.mit.edu/
VOC 0712 http://host.robots.ox.ac.uk/pascal/VOC/
VisDrone2019-DET https://github.com/VisDrone/VisDrone-Dataset
ExDark https://github.com/cs-chan/Exclusively-Dark-Image-Dataset
FLIR-ADAS2 https://www.flir.com/oem/adas/adas-dataset-form/
MMPretrain https://github.com/open-mmlab/mmpretrain
MMDetection https://github.com/open-mmlab/mmdetection
MMSegmentation https://github.com/open-mmlab/mmsegmentation
Appendix C Visualization of Detection Results
We visualize the detection results using detectors such as Faster R-CNN, Cascade R-CNN, and RetinaNet. For each detector, we randomly select two samples from MSCOCO val2017 and conduct comparisons with the ResNet-50 baseline. As shown in Figure 7, our method demonstrates superior performance in challenging scenarios, including obstruction, dense environments, clusters of small objects, and low-light conditions.
Appendix D More Experiments
@@
We are keen to explore whether attention mechanisms can be more effectively applied to various complex scene tasks. While previous research has shown good performance on general large-scale datasets, the effectiveness in dense, low-light, and small-object scenes remains uncharted. Therefore, we conducted more experiments using representative datasets: the small-object dataset VisDrone2019[48], low-light dataset ExDark[49], infrared automotive dataset FLIR-ADAS2[50], and general dataset VOC 0712[51]. The experimental setup was consistent with the detection configurations in Section IV. As shown in Table VI, it is gratifying that our SCSA outperformed others across these datasets, further demonstrating the robustness of our strategy in maintaining channel counts and the synergistic concept of multi-semantic information. Furthermore, our results indicate that the application of attention mechanisms on long-tail datasets, such as FLIR-ADAS2, has led to minimal performance gains and even declines. This may be due to the attention mechanism’s squeeze-and-excitation strategy being ill-suited for handling imbalanced distributed data, resulting in a focus on high-frequency categories while neglecting the learning of low-frequency ones. We believe that our approach will also perform well in other detection and segmentation tasks.
Appendix E Further Analysis of Effectiveness
Due to inherent semantic disparities across various objects and scales, our approach leverages these multi-semantic disparities. We employ shareable depth-wise convolutions to independently learn unique patterns in different features, effectively capturing the distinctiveness of each object. Additionally, by utilizing spatial knowledge that can perceive fine-grained and coarse-grained features of different objects, our method guides strong feature interactions and recalibration. This process integrates and alleviates the semantic ambiguities caused by semantic disparities, motivating our synergistic idea. For objects of different categories or the same category at different scales, the SMSA module extracts multi-semantic spatial knowledge, while the PCSA module efficiently integrates multi-scale semantic information. This enriches the global context, providing more comprehensive support for local decisions. Precise handling of local details helps selectively emphasize pixel-level areas of interest, which is a key factor in our significant advantages in object detection, instance segmentation, and semantic segmentation tasks.
E-A Visualization of ERFs
As depicted in Figure 8, leveraging the spatial structure of multi-semantic modeling, our SCSA has achieved a broader perceptual area. A larger effective receptive field(ERF) is beneficial for the network to utilize rich contextual information for collective decision-making, which is one of the important factors for performance improvement. To verify that the performance of our method benefits from a larger ERF, we randomly sample 300 images of different categories from the ImageNet-1K validation set, measure the contribution of each pixel on the original image to the center point of the output feature maps of the third and fourth stages of the model, and quantify the range of the ERF with the gradient values weighted and normalized. The visualization results demonstrate that as the network layers deepen, the ERF of our SCSA becomes increasingly evident, confirming our hypothesis and the effectiveness of our method.

### v2

@@ 3.3 Synergistic Effects
Baseline SCSA-50 2019 77.49
Macro Design w/o SMSA 2217 77.21
@@
Ordering PCSA Prior 2005 77.20
GN Prior 2010 77.47
Micro Design w/o Normalization in SMSA 2010 77.15
[…]
@@
Table 1: Ablation studies on the design strategy of SCSA, conducted at a 224x224 resolution, using the ImageNet-1K validation set. The abbreviation "PC" denotes progressive compression. $Gi(K_{1},K_{2},\ldots,K_{i})$ denotes splitting $X$ into $i$ sub-features and applying a 1D convolution of size $K_{i}$ to each $i$-th sub-feature.
Figure 3: Main Module Structures with SCSA
3.4 Integration of Attention Mechanisms
In our work, we integrate the proposed SCSA into different backbone networks to validate its effectiveness in enhancing feature extraction capabilities. As shown in fig. 3, the SCSA is integrated into four mainstream blocks: (a) and (b) represent blocks based on the ResNet [19] and its variant series [60]; (c) represents the inverted residual structure based on the MobileNet series [46, 21, 37]; (d) represents the block structure of RepVGG [8], a representative of the reparameterization approach.
@@ 3.4 Integration of Attention Mechanisms
ResNet-50 ResNet 25.56 4.02 2433 76.39 93.09
Channel ECANet 25.56 4.11 2109 77.05 93.43
SENet 28.07 4.11 2077 77.23 93.56
FcaNet 28.07 4.11 1905 77.29 93.64
Hybrid CPCA 27.40 4.87 1379 75.80 92.57
CBAM 28.07 4.12 1687 77.12 93.50
SANet 25.56 4.12 1493 77.12 93.64
ELA 25.59 4.11 2233 77.25 93.52
CA 25.69 4.11 2244 77.37 93.52
EMA 25.57 4.18 1861 77.43 93.79
@@
ResNet-101 ResNet 44.55 7.83 1588 77.76 93.81
Channel ECANet 44.55 7.83 1408 78.32 93.99
SENet 49.30 7.84 1399 78.40 94.05
FcaNet 49.29 7.84 1242 78.51 94.10
Hybrid CBAM 49.30 7.84 1118 78.09 94.07
CA 44.80 7.84 1437 78.11 93.92
@@
MobileNetV2-1.0 MobileNetV2 3.51 0.31 6693 71.54 90.11
Channel ECANet 3.51 0.31 5746 72.02 90.35
Hybrid CBAM 4.07 0.32 4539 72.43 90.49
@@
RepVGG-A0 RepVGG 9.11 1.52 6685 72.30 90.49
Channel ECANet 9.11 1.52 5059 72.76 90.71
Hybrid CA 9.35 1.52 3095 73.12 90.99
SCSA(Ours) 9.18 1.53 2842 73.51 91.12
Swin-T Swin 28.29 4.51 1523 80.83 95.49
Hybrid SCSA(Ours) 28.36 4.52 1315 81.53 95.84
Table 2: Comparison of our proposed SCSA with other state-of-the-art attention mechanisms across multiple benchmark models at a 224x224 resolution on the ImageNet-1K validation set [44].
4 Experiments
4.1 Experiments Settings
@@
Table 3: Comparison of the performance of different attention mechanisms for object detection on the MSCOCO validation set [30], utilizing models such as Faster R-CNN [43], Cascade R-CNN [3], and RetinaNet [29]. All models were fine-tuned using the "1×1$\times$1 ×" schedule with optimal results in bold and suboptimal results underlined in blue.
@@
In this section, we first introduce the experimental details. Next, we conduct experiments on four visual tasks, comparing our proposed SCSA with other state-of-the-art attention mechanisms. Following this, in section 4.5, we perform a comprehensive ablation study on our meticulously designed SCSA from four different perspectives.
Datasets.
We validate the effectiveness of our method across four visual tasks. For the image classification, we select the widely used ImageNet-1K [44] dataset. In the object detection, we employ several challenging detection datasets, including MSCOCO [30], Pascal VOC [11], VisDrone [10], and ExDark [34]. For semantic segmentation and instance segmentation, we selected the widely used ADE20K [67] and MSCOCO [30] benchmarks.
We are keen to explore whether attention mechanisms can be more effectively applied to various complex scene tasks. While previous attention research [23, 55, 65, 39, 57, 61] has shown good performance on widely used benchmarks (e.g., ImageNet-1K [44], MSCOCO [30]), the effectiveness in dense, low-light, and small-object scenes remains uncharted. Therefore, we conduct more experiments using representative benchmarks in table 4: the small-object dataset VisDrone [10], low-light dataset ExDark [34], infrared automotive dataset FLIR-ADAS v2 [13], and general dataset Pascal VOC [11].
Metrics.
We use Top-1 and Top-5 metrics to measure image classification, Average Precision (AP) to evaluate object detection, and report Parameter Count (Params) and Floating Point Operations Per Second (FLOPs), and throughput to measure performance. For semantic segmentation, we employ the mean Intersection over Union (mIoU).
Implementation Details.
@@
To evaluate our proposed SCSA on ImageNet-1K [44], we select four mainstream backbone networks based on CNN and Transformer architectures, including ResNet [19], MobileNetV2 [46], RepVGG [8] and Swin [33]. Specifically, we follow the parameter configurations in the original papers [19, 46, 8, 33], except for the batch size and learning rate. Since all classification models are trained on a single NVIDIA RTX 4090 GPU, we adjust the batch size and learning rate according to the linear scaling rule [15, 63]. For ResNet [19], RepVGG [8], and Swin [33], the batch size is uniformly set to 128, with the learning rates scaled down to 0.05, 0.05, and 0.000125, respectively. When training MobileNetV2 with our SCSA, we use the batch size and learning rate from ECANet [55], set to 96 and 0.045, respectively. Notably, to enhance training efficiency, we employ Automatically Mixed Precision(AMP) training.
@@
We evaluate our SCSA on MSCOCO [30] using Faster R-CNN [43], Mask R-CNN [18], Cascade R-CNN [3], and RetinaNet [29]. These detectors are implemented using the MMDetection [5] toolboxes with default settings. All models are trained using an SGD optimizer with a momentum of 0.9 and a weight decay of 1e-4, with a batch size of 2 per GPU, over a total of 12 epochs. Faster R-CNN, Mask R-CNN and Cascade R-CNN started with a learning rate of 0.0025, while RetinaNet starts at 0.00125. The learning rates for all models are decreased by a factor of 10 at the 8th and 11th epochs. We fine-tuned the model on MSCOCO [30] for 12 epochs using a single NVIDIA H800 GPU and reported comparative results on validation set. Building on the above configurations, we further evaluate the proposed SCSA method’s detection performance and generalization capability on Pascal VOC [11], as well as in complex scenarios such as VisDrone [10], ExDark [34], and FLIR-ADAS V2 [13].
@@
We further validate our method on ADE20K [67] with the UperNet [59] for semantic segmentation. Following common practices [6, 64], we utilize the MMSegmentation [7] toolboxes, set the batch size to 16, and conduct 80k training iterations. All models are trained using an SGD optimizer with an initial learning rate of 0.01, a momentum of 0.9, and a weight decay of 5e-4. We also conduct training and inference using a single NVIDIA H800 GPU.
@@ 4.2 Image Classification
We compare our SCSA against other state-of-the-art attention mechanisms, including SENet [23], CBAM [57], ECANet [55], FcaNet (FCA) [41], CA [20], SANet [65], EMA [39], CPCA [24] and ELA [61]. As shown in table 2, our SCSA achieved the highest Top-1 accuracy across networks of different scales, with negligible parameter count and computational complexity. Within hybrid architectures, our method’s throughput based on ResNet is second only to CA and ELA, but it offer a better balance of accuracy, speed, and model complexity with a moderate model width. Integrating the SCSA method into the MobileNetV2 architecture significantly improves model accuracy. Although SCSA is more lightweight in terms of parameter count (3.63M vs. 4.07M, -0.44M), its multi-branch structure encounters a sharp increase in channel dimensions within the inverted residual blocks, which leads to a reduction in throughput. Notably, integrating the proposed SCSA method into advanced models like RepVGG [8] and the spatial self-attention-based Swin [33] still yields notable accuracy improvements of 1.21% and 0.70%, respectively, effectively demonstrating the adaptability of our attention mechanism across different model architectures.
4.3 Object Detection
Results on MSCOCO.
We evaluate various attention mechanisms on MSCOCO to verify the effectiveness of our approach in dense detection scenario. We use ResNet-50 and ResNet-101 as the backbone and FPN [28] as the feature fusion network. As shown in table 3, our method outperforms other state-of-the-art attention methods across various detectors, model sizes, and object scales. For Faster R-CNN [43], our SCSA improves by 1.7% and 1.3% in terms of AP compared to the original ResNet-50 and ResNet-101, respectively. Compared to other plug-and-play attention modules, including CBAM [57], FCA [57], ECA [55], and CA [20], SCSA demonstrates superior performance, achieving gains of 0.4% to 1.0% on the Cascade R-CNN [3] detector. Moreover, it consistently excels in detecting targets across a range of scales, confirming its robust adaptability to multi-scale features.
@@ 4.3 Object Detection
Table 4: Comparison of our SCSA, based on ResNet-50 and ResNet-101, with other attention mechanisms for object detection performance across four different datasets.
Results on Infrared, Low-Light, and Small Target Detection.
@@
As shown in table 4, it is gratifying to see that proposed SCSA performs better across these benchmarks [34, 10, 11, 13] compared to other counterparts, further demonstrating the robustness of our strategy in maintaining channel dimensions and the synergistic concept of multi-semantic information. Notably, our results indicate that there are still some limitations in the application of attention mechanisms on long-tail datasets, such as FLIR-ADASv2 [13], has led to minimal performance gains and even declines. This may be due to the attention mechanism’s squeeze-and-excitation strategy being illsuited for handling imbalanced distributed data, resulting in a focus on high-frequency categories while neglecting the learning of low-frequency ones.
Methods UperNet
Params(M) FLOPs(G) mIoU(%)
ResNet-50 64.10 1895 40.20
+ CBAM 66.62 1895 39.62
+ CPCA 65.94 1927 39.68
+ SE 66.62 1895 39.94
+ SA 64.10 1895 40.01
+ ECA 64.10 1895 40.46
+ FCA 66.61 1895 41.09
+ SCSA(Ours) 64.16 1895 41.14
ResNet-101 83.09 2051 42.74
+ CBAM 87.84 2051 41.65
+ ECA 83.09 2051 42.63
+ SE 87.84 2051 42.66
+ FCA 87.83 2051 43.22
+ SCSA(Ours) 83.22 2051 43.76
Table 5: Comparison of our method, based on the UperNet model, with other attention mechanisms for semantic segmentation performance on the ADE20K benchmark.
Methods Mask R-CNN
AP $AP_{50}$ $AP_{75}$ $AP_{S}$ $AP_{M}$ $AP_{L}$
ResNet50 34.8 55.9 36.9 16.4 37.4 50.2
+ CBAM 35.4 56.9 37.6 17.4 38.3 50.6
+ ECA 35.5 57.6 37.6 16.6 38.4 52.0
+ FCA 35.5 57.2 37.6 17.1 38.6 51.3
+ SE 35.7 57.3 38.1 17.7 38.6 50.9
+ SA 35.7 57.7 38.0 17.2 38.7 51.5
+ CA 35.8 57.5 38.2 16.9 38.5 51.7
+ SCSA 36.1 58.4 38.3 17.2 39.1 51.9
Table 6: Comparison of our method, based on the Mask R-CNN, with other attention mechanisms for instance segmentation performance on the MSCOCO validation set [30].
4.4 Segmentation
We also test its performance in semantic segmentation on ADE20K [67] and instance segmentation on MSCOCO [30]. We conduct extensive comparative experiments based on the UperNet [59] network. As shown in table 5 and table 6, our SCSA significantly outperforms other attention methods. Specifically, SCSA improves performance by 0.94% and 1.02% in terms of mIoU on ResNet-50 and ResNet-101, respectively, while other methods only achieve improvements of 0.1% to 0.2%, and some even fall below the baseline model. Meanwhile, SCSA achieves a 0.3% to 0.7% increase in terms of AP in instance segmentation tasks, surpassing other counterparts. These results demonstrate that our method, based on multi-semantic spatial information, performs well in pixel-level tasks.
4.5 Ablation study
As shown in table 1, we apply SCSA to ResNet-50, constructing SCSA-50 as the baseline on ImageNet-1K [44] for ablation studies across four aspects.
Macro Design.
We validate the SMSA and PCSA modules separately, and both show significant improvements in accuracy compared to ResNet-50. With SMSA, guided by multi-semantic information, the Top-1 accuracy improved significantly by 1.05%, while PCSA, which mitigates multi-semantic disparities and promotes channel interaction, increased the accuracy by 0.82%. Without progressive compression in PCSA, accuracy drops by 0.18%, and this is primarily because, after direct global spatial compression, PCSA cannot leverage the discriminative spatial priors provided by SMSA for its computations.
Ordering.
Our study primarily aims to explore whether the inherent multi-semantic spatial information across the spatial and channel dimensions can effectively guide the learning of channel features. To further demonstrate the benefits of this "guidance," we swapped the order of PCSA and SMSA. Interestingly, the Top-1 accuracy dropped by 0.29%, which further validates our previous hypothesis that spatial attention can guide channel feature learning, thereby confirming the effectiveness of guiding with multi-semantic information.
Although normalization helps reduce the impact of data noise and accelerates model convergence [45], the placement of normalization within SMSA may yield varying effects. Backbone networks based on MHSA [52, 9] typically use Layer Normalization (LN) [1] before attention computation, whereas some plug-and-play attention modules either omit normalization layers [55, 24] or apply them beforehand [65, 39]. To explore the necessity and optimal placement of normalization in SMSA, we conduct experiments by placing normalization before attention and by removing it. Results shown in table 1 indicate that normalization is essential for the attention mechanism, though its specific position has a minor impact. Pre-normalization aids in handling variations among input features and improves training stability, but may cause loss of feature details, reducing attention sensitivity to fine-grained information. In contrast, applying normalization after attention calculation can mitigate noise but may also diminish the model’s focus on important features. Ultimately, based on accuracy results, we opted to place normalization after attention calculation in SMSA.
Micro Design.
Following the experimental analysis above, which confirms the importance of normalization layers in attention calculation, we consider whether group normalization is more suitable for extracting multi-semantic information from multiple sub-features in the proposed SCSA. To investigate this, we perform ablation studies comparing popular normalization methods in deep neural networks (DNNs), such as BN [26] and LN [1]. The results show that when GN is sequentially replaced with BN and LN, both accuracy and inference speed decrease, with Top-1 accuracy dropping to 77.19% and 77.20%, respectively. These declines are attributed to GN’s superior ability to preserve the independence of semantic patterns among sub-features, thereby minimizing semantic interference. Conversely, BN’s sensitivity to batch size can introduce statistical noise [2] when processing multi-semantic information. LN, by normalizing along the channel dimension and capturing information across all features, may disrupt the distinct semantic patterns that SMSA’s multi-scale convolutions extract from individual sub-features. These ablation suggest that GN may be a more suitable choice in convolution layers that involve multiple semantics. Furthermore, the decline in accuracy and increase in parameters with unshared convolutions further validate the effectiveness of using shared convolutions to consistently learn and model features dependencies across the H and W dimensions.
Additionally, when replacing the single attention head in PCSA with multi-head and channel shuffle operation [66], performance decrease from 77.49% to 77.35%. This phenomenon is primarily attributed to the strong inter-channel interactions facilitated by the single head, which effectively alleviate semantic disparities produced in SMSA. To validate the shared convolutional learning on 1D sequences decomposed along the Height and Width dimensions, we compare it to non-shared convolutional learning. Results show a 0.17% accuracy drop, reduced throughput, and increased parameters and FLOPs due to more convolutional operators. This confirms that shared learning across dimensions captures complementary features, enhancing model expressiveness.